Suppose you are given the following closed economy model:


C = 200 + 0.78Yd


T = 0.18Y


I = 50


G = 800



i. Calculate the equilibrium level of income and consumption. (5 Marks)


ii. Calculate the government expenditure multiplier and interpret the result. (4 Marks)


iii. Find the fiscal deficit/surplus and interpret the result. (3 Marks)


iv. Suppose government purchases increase to 1000, what is the new equilibrium income and consumption? (4 Marks)


v. What is the new fiscal deficit/surplus after government spending increases to 1000?



1
Expert's answer
2022-03-09T11:03:15-0500

Solution

i)Calculate the equilibrium level of income and consumption.

The expenditure approach method is given by;

"Y=C+I+G+(X-M)"

Therefore;

"Y=200+0.78(Y-0.18Y)+50+800"

"Y=200+0.78Y-0.1404+50+800"

Put like terms together and solve

"Y-0.78Y=200-0.1404+50+800"

"0.22Y=1049.8596"

"Y=\frac{1049.8596}{0.22}"

"Y=4,772.10"

"Y=" $4, 772.10


ii)Calculate the government expenditure multiplier and interpret the result.

"G=\frac{change in total spending}{change in autonomous spending}"

"G=\frac{4772.10}{800}"

"G=" 5.97



iii)Surplus or deficit

surplus/deficity = cash receipts-cash disbursements

"=47772.10-(859+200+859)"

"=" 2,854.10

There's a surplus of $2,854.10 in the economy.